i'm not quite sure how to solve this using substitution.
3x-2y=11
x-y=4
Also, does the system have exactly one solution, or more?
Thanks for helping me!

Answer by Abbey(339)

3x-2y=11
x-y=4
add y to both sides of the second equation to get
x=y+4
then use Y+4 in the first equation instead of x
3(y+4)-2y=11
3y+12-2y=11
y+12=11
y=-1
then put -1 back into your equation and solve for x:
x-y=4
x-(-1)=4
x+1=4
x=3
the solution is (3,-1)
This will have only one solution because these are two straight lines, and the solution is where these two lines cross. (straight lines can only cross once).